OpenStudy (anonymous):

y= √ x (x+1) derivatives

OpenStudy (anonymous):

Use the chain rule\[\frac{d}{dx} \sqrt{x(x+1)}=\frac{2x+1}{2 \sqrt{x(x+1)}}\]

OpenStudy (anonymous):

1 +1/2x^1/2 would be answer right

OpenStudy (anonymous):

If this is your function: \[\sqrt{x(x+1)}\] the derivative is \[\frac{2x+1}{2 \sqrt{x(x+1)}}\] Which is your answer

OpenStudy (anonymous):

okay thanks